To make a solution, 38.7 g of a nonelectrolyte is dissolved into 218 g of water. The freezing point of the solution is measured to be -5.53°C. Calculate the molar mass of the solute.

1
Expert's answer
2022-03-12T01:58:37-0500

∆T= iKfm

For water, Kf=1.86oc/kg/mol

For a non-electrolyte, i=1 since it does not dissociate into ions.

m=molality of solution

m=(38.7×1000)÷(218×molar mass)

= 177.5229÷molar mass


5.53=1×1.86×177.5229÷molar mass

Molar mass= 1.86×177.5229÷5.53

=59.7